# An Innovative Attention-based Ensemble System for Credit Card Fraud Detection

### v1

@@ Abstract
Detecting credit card fraud (CCF) holds significant importance due to its role in safeguarding consumers from unauthorized transactions that have the potential to result in financial detriment and negative impacts on their credit rating. It aids financial institutions in upholding the reliability of their payment mechanisms and circumventing the expensive procedure of compensating for deceitful transactions. The utilization of Artificial Intelligence methodologies demonstrated remarkable efficacy in the identification of credit card fraud instances. Within this study, we present a unique attention-based ensemble model. This model is enhanced by adding an attention layer for integration of first layer classifiers’ predictions and a selection layer for choosing the best integrated value. The attention layer is implemented with two aggregation operators: dependent ordered weighted averaging (DOWA) and induced ordered weighted averaging (IOWA). The performance of the IOWA operator is very close to the learning algorithm in neural networks which is based on the gradient descent optimization method, and performing the DOWA operator is based on weakening the classifiers that make outlier predictions compared to other learners. Both operators have a sufficient level of complexity for the recognition of complex patterns. Accuracy and diversity are the two criteria we use for selecting the classifiers whose predictions are to be integrated by the two aggregation operators. Using a bootstrap forest, we identify the 13 most significant features of the dataset that contribute the most to CCF detection and use them to feed the proposed model. Exhibiting its efficacy, the ensemble model attains an accuracy of 99.95% with an area under the curve (AUC) of 1.
@@
Credit card fraud detection, Fusion-based parallel ensemble system, Attention layer, Machine learning, Classifier selection, Aggregation operators.
@@ I Introduction
The convenience of financial services for consumers around the world has been tremendously improved by digital technology. However, this development also opened up new avenues for financial fraud, especially in credit card transactions [1, 2], [3]. Criminals defraud people through various technologies and false identities intending to steal their money [4]. This gave way to a steep increase in credit card frauds as e-commerce and online shopping gained momentum [5, 6].
@@
Looking at the numbers, one can tell that credit card fraud (CCF) is a big concern. According to Nilson’s report; global losses from card fraud amounted to 28.65 billion dollars in 2019. These figures are expected surpassing 43 billion dollars by 2025 [7, 8]. The large increase shows that fraudulent methods are getting more complicated day by day, hence making it hard to beat fraudsters at their game [9, 10]. In the United States alone, CCF cases reported to the Federal Trade Commission (FTC) were over 393 thousand during 2020, indicating a significant rise from the previous year [11, 12]. The advent of internet-based transactions worsened matters leading into rise of fraud cases involving card-not-present (CNP) transactions, which are quite common currently [1, 13]. Compared to point-of-sale fraud, the prevalence of CNP fraud rose by 81%, suggesting a major shift in digital crime [14]. This therefore means that it is necessary to propose models and techniques for detecting deceitful transactions. The old ways of identifying fraud were not quick enough [15, 16]. As a result, researchers performed thorough investigations and practically used machine learning (ML) algorithms specifically designed for CCF detection [17, 18, 19, 20].
Fraud detection extensively relies on the utilization of ML algorithms. Various ML algorithms, like support vector machine (SVM), decision tree, random forest, LightGBM, XGBoost, naïve bayes, stacking classifier and voting classifier are utilized in fraud detection [21, 22, 23, 21, 24].
Deep learning (DL) techniques were widely used in detecting fraud across fields in CCF detection. Models like Convolutional Neural Networks (CNN), Recurrent Neural Networks, and Autoencoders displayed results in spotting activities [25, 1, 26]. Hybrid models such as CNN-SVM, Autoencoder-LightGBM, ensemble method and others were commonly employed for CCF detection [27, 28, 29, 30].
Our key contributions to this research are:
  • We introduce a novel fusion-based ensemble system by adding an attention layer and a selection module to the traditional stacking classifier.
  • Intelligently, two aggregation operators, DOWA and IOWA, are employed to combine the predictions of first layer classifiers. In this entirely fusion-based architecture, a synergistic combination of the aggregated values by these two operators is used to feed the meta-learner.
  • Dimension reduction using the bootstrap forest approach is accomplished by eliminating features that contribute negligibly to the prediction accuracy.
  • The best classifiers are selected for combination based on two criteria: performance and diversity.
  • The proposed model demonstrates 99.95% accuracy in detecting credit card fraud with the AUC of 1.
The rest of this paper is structured as follows: Section II introduces related work on detecting CCF. Section III first addresses the dataset and the identification of its most significant features, then introduces the aggregation operators used in the attention layer, and finally analyses the structure of the proposed system. Findings and analysis are examined in Section IV. Section V presents prospects for future studies. Finally, Section VI considers the conclusion.
II Background and Related Work
Several methodologies and algorithms were deployed to predict fraud, which saw a sharp rise in recent times. A focal area of various studies is the detection of CCF. These paragraphs delve into some of these research efforts.
Long et al. A graph network model was introduced to efficiently detect online fraud by addressing challenges such as feature disparities, category imbalances, and relationship issues [31, 32, 33, 34].
In [12], the authors proposed a CCF detection model using new transactional behavioral representations. Two time-aware gates were designed in a recurrent neural network to capture long- and short-term transactional habits and behavioral changes influenced by time intervals between transactions. The model used a time-aware attention module to extract behavioral information from historical transactions, considering time intervals, and an interaction module to enhance the learned representations. Similarly, in the realm of spotting news, LSTM and Bi LSTM methods were utilized for text data preprocessing and analysis, achieving a level of accuracy in detecting misinformation [35, 36].
Fidelis and coworkers in [37] presented a novel approach using a hybrid cultural genetic algorithm (CGA) alongside a trained modular neural network ensemble for classifying transactions. This ensemble included a complex architecture of modular neural networks, featuring components such as an encoder, selectors, a swapped recombine, swapped dynamic amplification, and a belief finisher for CGA operations. The ensemble relied on optimized rules derived from the GA-block for effective classification tasks.
Ibomoiye et al. in [38] introduced a DL ensemble method to identify CCF utilizing a mix of LSTM and gated recurrent unit (GRU) networks, as primary learners alongside an MLP acting as the meta-learner. The combined model demonstrated superior performance in sensitivity, specificity and AUC metrics when contrasted with performing GRU, LSTM and MLP models.
Within [39], the authors presented an innovative approach to detecting fraud, comprising four distinct stages. Initially, cardholders were classified into groups based on similar transaction behaviors using historical transaction data. Next, a window-sliding strategy is utilized to merge transactions in each group. Following that, distinct behavioral patterns are derived from every cardholder. Separate classifiers are trained for each group, enabling real-time fraud detection.
In [40], the authors propose the utilization of an adaptive sampling and aggregation-based graph neural network (ASA-GNN) in order to enhance transaction fraud detection. By employing a neighbor sampling strategy that selectively filters out noisy nodes and provides supplementary information for fraudulent nodes, the model effectively addresses the challenge of learning discriminative representations. To selectively choose similar neighbors with comparable behavior patterns, cosine similarity and edge weights are utilized, while introducing a neighbor diversity metric to address fraudsters’ attempts at disguise and prevent excessive smoothing.
In [41], Wang and colleagues introduced an ensemble approach that incorporates various feature ranking methods to achieve more reliable results. This method involved a two-step process: first, employing multiple rankers to create different feature rankings, and then combining these rankings using a median-based method to determine the top ten features for model building. This approach led to the development of models utilizing Decision Tree-based classifiers, specifically CatBoost and XGBoost, enhancing predictive stability and dependability.
Esenogho et al. described an ensemble classifier combining an LSTM network with the AdaBoost algorithm for efficient credit card fraud detection in [42]. Here, the LSTM serves as the base learner within AdaBoost, using data refined by the synthetic minority over-sampling technique-edited nearest neighbor (SMOTE-ENN) technique for better model training. The ensemble’s final decision was reached through a weighted voting scheme among the outcomes of various LSTM models, aiming to boost fraud detection accuracy.
In [43], the authors introduced a fresh approach to detect consumer fraud in reviews. Their method involves integrating a channel biattention convolutional neural network (CNN) with a pre-trained language model. A novel similarity computation module is introduced, which employs a metric matrix to evaluate the connection between prior knowledge and consumer reviews. This enables the model to effectively identify previously undetected fraudulent behaviors.
Although various efforts were made towards the detection and prediction of CCF, many focused on single classifiers, which are not as reliable as ensemble-based classification methods. This is because multi-classifier approaches use the consensus of several single classifiers, resulting in a more reliable outcome. Studies that examined CCF detection using ensemble classifiers either employed outdated versions such as bagging, boosting, and stacking.
Figure 1: Heatmap of correlation matrix.
III Proposed Attention-based Ensemble System for CCF Detection
In this section, a new parallel ensemble architecture based on information fusion is introduced. This novel architecture is constructed by adding two layers to the stacking classifier structure: a) an attention layer, and b) a selection layer. Information fusion in the attention layer is accomplished through two operators, DOWA and IOWA, while the selection layer employs a mathematical logic to choose one of the two output vectors from the attention layer for each sample. Ultimately, the selected vector feeds the meta-learner.
III-A Dataset
The dataset in question includes over 550,000 anonymized records of credit card transactions performed by European cardholders in 2023 [44]. The anonymization was meticulously done to ensure the privacy and security of the cardholders’ identities. This dataset’s main purpose is to aid in creating and honing fraud detection algorithms and systems capable of pinpointing transactions that might be fraudulent. Some principal characteristics of this dataset are:
  • The dataset comprises a unique identifier, named “id” assigned to each transaction for distinct identification.
  • It includes anonymized features labeled V1 through V28 that encapsulate varying aspects of the transactions, such as timing and geographical details, among others, in a manner protective of user privacy.
  • “Amount” denotes the monetary value of the transaction.
  • there is a binary attribute named “Class” which serves the purpose of classifying transactions into two categories: fraudulent (marked as 1) and legitimate (marked as 0).
Figure 2: this boxplot visualizes the distribution of dataset features along with outliers.
Fig. 1 shows the correlation matrix, which displays the extent of linear dependencies among different features of the dataset through a heatmap. The more a cell corresponding to two features tends towards violet or white, it indicates higher linear dependencies. The boxplot in Fig. 2 Offers a graphical representation of the essential characteristics of the dataset, encompassing the median, upper and lower quartiles, and potential outliers.
Figure 3: Predictor screening report for feature selection
III-B Predictor screening
The process of predictor screening is vital for singling out significant predictors from a broad pool of potential ones. This technique uses a bootstrap forest method to identify predictors that could impact the outcome variable.
In the analysis, a set of 100 decision trees is created using the bootstrap forest approach for each response. This model evaluates how each predictor influences the prediction of outcome. The predictors are ranked based on their contributions to the model. Fig. 3 illustrates this ranking, displaying the predictors with their corresponding contributions.
Hence, predictors whose contribution percentage falls below 0.5% are classified as insignificant dimensions. By removing these negligible dimensions from feature space $R^{29}$, we effectively transition into feature space $R^{13}$. Thus, as per Fig. 1 and Fig. 3, 16 features that have high linear or non-linear dependency on each other or are not sufficiently discriminative are removed from the raw dataset to prevent the model from becoming biased towards these uninformative features.
III-C aggregation operators
The OWA operator is a mathematical tool used for merging inputs by considering both their order and assigned weights, frequently applied in processes of decision-making and aggregation. This method systematically orders the inputs before applying specific weights to them, as highlighted in [45]. Definition 1 introduces the DOWA operator, and the IOWA operator is introduced in definition 2. These two operators are used separately to combine the predictions of first layer classifiers in subsection III-D.
@@ III-B Predictor screening
$w_{j}=\frac{c(p_{\alpha(j)},m)}{\sum\limits_{j=1}^{n}c(p_{j},m)}.\ \ \ \ j=1,2,...,n,$ (4)
@@
We are presented with a set of samples, each comprising a series of values $(p_{k1},p_{k2},...,p_{kn})$, referred to as the arguments, and a corresponding single value named the aggregated value, denoted as $d_{k}$. Our aim is to develop a model that accurately captures this aggregation process using the OWA (Ordered Weighted Averaging) method. In essence, the task is to derive a weighting vector $W$ that effectively represents the aggregation mechanism across the dataset. The aim is to determine the weighting vector $W$ based on the data provided.
@@
Minimize the immediate errors $\textbf{e}_{k}:$
@@
$\beta_{i}(l+1)=\beta_{i}(l)-\alpha w_{i}(b_{ki}-\hat{d}_{k})(\hat{d}_{k}-d_{k})$ (12)
@@
$w_{i}=\frac{e^{\beta_{i}(l)}}{\sum\limits_{j=1}^{n}e^{\beta_{j}(l)}},\ \ i=[1,n].$ (13)
@@
Using the 10-fold cross-validation method, we divide the predicted probabilities into two groups: test data and training data. Then, with the help of the proposed algorithm, we learn the weights from the training data, and then, using these weights and the sorted input arguments, the aggregation value is calculated for the test data.
III-D CCF detection Framework: Structure, Development, and Examination
@@
Fig. 4 illustrates the structure of the traditional stacking classifier. Stacking is a method of ensemble learning that merges predictions from base models to generate a final prediction.
These base models can be of various types, including decision trees, SVM, neural networks, etc. The predictions made by these base models are then used as input features for a second-level model (or meta-learner) that aims to correct the predictions of the base models and produce the final prediction.

### v2

# Explainable AI for Fraud Detection: An Attention-Based Ensemble of CNNs, GNNs, and A Confidence-Driven Gating Mechanism

@@ Abstract
The rapid expansion of e-commerce and the widespread use of credit cards in online purchases and financial transactions have significantly heightened the importance of promptly and accurately detecting credit card fraud (CCF). Not only do fraudulent activities in financial transactions lead to substantial monetary losses for banks and financial institutions, but they also undermine user trust in digital services. This study presents a new stacking-based approach for CCF detection by adding two extra layers to the usual classification process: an attention layer and a confidence-based combination layer. In the attention layer, we combine soft outputs from a convolutional neural network (CNN) and a recurrent neural network (RNN) using the dependent ordered weighted averaging (DOWA) operator, and from a graph neural network (GNN) and a long short-term memory (LSTM) network using the induced ordered weighted averaging (IOWA) operator. These weighted outputs capture different predictive signals, increasing the model’s accuracy. Next, in the confidence-based layer, we select whichever aggregate (DOWA or IOWA) shows lower uncertainty to feed into a meta-learner. To make the model more explainable, we use shapley additive explanations (SHAP) to identify the top ten most important features for distinguishing between fraud and normal transactions. These features are then used in our attention-based model. Experiments on three datasets show that our method achieves high accuracy and robust generalization, making it effective for CCF detection.
@@
Credit card fraud detection, Attention mechanism, Explainable AI, Information fusion, Ensemble learning.
@@ I Introduction
The convenience of financial services for consumers around the world has been tremendously improved by digital technology. However, this development also opened up new avenues for financial fraud, especially in credit card transactions [1, 2]. Looking at the numbers, one can tell that credit card fraud (CCF) is a big concern. According to Nilson’s report; global losses from card fraud amounted to 28.65 billion dollars in 2019. These figures are expected surpassing 38.5 billion dollars by 2027 [3, 4]. In the United States alone, CCF cases reported to the federal trade commission (FTC) were over 393 thousand during 2020, indicating a significant rise from the previous year [5, 6]. The advent of internet-based transactions worsened matters leading into rise of fraud cases involving card-not-present (CNP) transactions, which are quite common currently [1, 7].
Traditional fraud detection methods, such as rule-based systems and manual transaction reviews, are often inadequate in the face of evolving fraud patterns [8, 9]. These approaches are limited in their ability to adapt to new tactics employed by fraudsters and frequently generate high false-positive rates, leading to unnecessary inconvenience for legitimate customers [10, 11]. In contrast, machine learning (ML) and deep learning (DL) methodologies offer data-driven solutions that are capable of learning complex patterns and adapting to emerging fraud strategies [12, 13]. By analyzing vast amounts of transaction data, these techniques provide more accurate, scalable, and efficient ways to identify fraudulent transactions.
Recent progress in data-driven computational intelligence has introduced powerful classification algorithms, such as decision trees, random forests, gradient boosting, convolutional neural networks (CNNs), and recurrent neural networks (RNNs), specifically tailored for fraud detection [14, 15, 16, 17]. Additionally, ensemble techniques that synthesize the outputs of multiple models have shown notable success by exploiting the strengths of diverse classifiers [14, 18]. However, despite these achievements, critical challenges persist in balancing predictive accuracy ($A_{c}$) with interpretability, particularly within financial institutions that require clear, trustworthy insights into automated decisions [19].
Despite these notable advances, significant research gaps persist in CCF detection. First, many approaches still rely on a single classification model or a simple ensemble of models, which may fail to leverage complementary predictive signals and overlook more sophisticated aggregation strategies [20]. Second, although graph neural networks (GNNs), CNNs, and recurrent architectures like RNNs and long short-term memories (LSTMs) each show promise individually, their combined outputs are often integrated through naive averaging or straightforward stacking, diluting critical predictive cues. Third, most current work does not adequately address model uncertainty, potentially leading to unreliable predictions, particularly in highly imbalanced datasets. Finally, interpretability remains a pressing concern: despite its importance in promoting stakeholder trust, the majority of methods do not offer transparent or user-friendly ways to understand feature importance and decision rationale. These limitations underscore the need for a more robust, uncertainty-aware, and interpretable framework that intelligently fuses multiple neural models to enhance both $A_{c}$ and trust in credit card fraud detection systems. Our main contributions to this study are outlined below:
  • A diversity-based classifier selection is proposed, where an attention-based stacking framework is designed to leverage four distinct classifiers- CNN, RNN, LSTM, and GNN- to achieve a balanced trade-off between reliability and robustness. Specifically, CNN and RNN outputs are combined using the dependent ordered weighted averaging (DOWA) operator, while LSTM and GNN outputs are fused via the induced ordered weighted averaging (IOWA) operator.
  • A confidence-aware combination layer is introduced, enabling dynamic selection of the more reliable aggregate (DOWA or IOWA) based on uncertainty estimates. This best-performing output is then fed into a meta-learner, ensuring higher $A_{c}$, particularly under highly imbalanced conditions.
  • Interpretability is enhanced through the use of Shapley Additive Explanations (SHAP), allowing the identification of the top ten most influential features affecting model decisions. This feature-level insight improves transparency and facilitates stakeholder trust in the fraud detection process.
  • An extensive evaluation and benchmarking are conducted using three benchmark datasets, where the proposed method is compared against individual classifiers (CNN, RNN, LSTM, GNN). The results demonstrate superior performance and robust generalization, validating the effectiveness of the proposed framework in credit card fraud detection.
The remainder of this paper is organized as follows. Section II introduces the datasets used in our experiments and identifies the key features driving the proposed approach. It then discusses the aggregation operators (DOWA and IOWA) and concludes with a detailed exposition of the system’s architecture. Section III presents the experimental results, offering a comprehensive analysis of the findings. Section IV provides a broader discussion of these results, highlighting their implications and suggesting potential avenues for future research. Finally, Section V provides the concluding remarks.
II Proposed Attention-based Ensemble System for CCF Detection
In this section, we detail the methodology adopted to develop and evaluate our stacking-based approach for CCF detection. We begin by introducing the three benchmark datasets, highlighting their characteristics and the rationale behind their selection. We then describe the process for identifying key features using SHAP, focusing on the top attributes that distinguish fraudulent from legitimate transactions. Next, we explain how the DOWA and IOWA operators are employed to fuse the outputs of our neural network classifiers, CNN, RNN, LSTM, and GNN, in an attention mechanism. Finally, we outline the overall system architecture, including the confidence-based combination layer, which helps select the most reliable ensemble output before passing it to a meta-learner, ensuring both high $A_{c}$ and robust generalization.
II-A Datasets
For fraud detection, we use three datasets, each containing anonymized transaction-related features labeled $V1$ through $V28$ to preserve user privacy. These features capture various aspects of the transaction process, ranging from timing information to geographical context. Each dataset also includes the feature Amount, representing the monetary value of the transaction, and the binary attribute Class, where transactions are labeled as fraudulent (1) or legitimate (0). Table I provides a statistical overview of all three datasets. Notably, Dataset 3 was constructed by concatenating Dataset 1 and Dataset 2, resulting in a larger sample size for more comprehensive analysis and model training.
To rigorously assess our framework in both balanced and imbalanced scenarios, we first employ a balanced dataset (Dataset 1) for initial training and testing. This approach allows us to observe the model’s baseline performance when both classes are represented more evenly. Next, we train our proposed framework on the highly imbalanced Dataset 2 and subsequently evaluate it on this same dataset to investigate how the skew in class distribution influences predictive $A_{c}$. Finally, we tune the model’s hyperparameters on the balanced dataset and test the refined model on the unbalanced dataset, mirroring real-world conditions where fraudulent transactions constitute only a small fraction of all transactions. Since the features in two datasets were anonymized, we assumed their equivalence when training on Dataset 1 and testing on Dataset 2. This step-by-step procedure ensures that our model is both reliable under balanced conditions and robust to the challenges posed by heavily imbalanced data.
TABLE I: Data partition.
Datasets samples Fraud Imbalance ratio
Public 1 Train set 454 905 227 452 50%
Test set 113 725 56 863 50%
Public 2 Train set 227 845 394 0.17%
Test set 56 962 98 0.17%
Public 3 Train set 568 630 284 315 50%
Test set 284 807 492 0.17%
Figure 1: SHAP-based feature importance ranking for fraud detection.
II-B Feature Importance Analysis Using SHAP
We employ SHAP with a bootstrap-based approach using 100 decision trees to assess feature importance in our fraud detection model. The results, visualized in Fig. 1, rank features by their average SHAP values. The analysis highlights $V14$, $V10$, and $V4$ as the most influential features, while others, such as $V24$ and $V27$, have minimal impact. Interestingly, the $Amount$ feature contributes negligibly, indicating that fraud detection relies more on transaction behavior than monetary values. To improve efficiency and reduce complexity, we selected only the top 10 features ($V14$, $V10$, $V4$, $V17$, $V12$, $V11$, $V16$, $V8$, $V3$, and $V1$) for the final model. This selection enhances interpretability, reduces overfitting risks, and ensures practical deployment while maintaining high predictive performance.
II-C Aggregation Operators, DOWA and IOWA
The Ordered Weighted Averaging (OWA) operator is a mathematical tool commonly employed in decision-making and aggregation tasks. It arranges inputs in a specified order before applying designated weights, as highlighted in [21]. In this study, Definition 1 introduces the DOWA operator, while Definition 2 presents the IOWA operator. In section II-D, each operator is used independently to merge the predictions of the first-layer classifiers within our proposed framework.
@@ II-C Aggregation Operators, DOWA and IOWA
$w_{j}=\frac{c(p_{\alpha(j)},m)}{\sum\limits_{j=1}^{n}c(p_{j},m)},\ \ \ \ j=1,2,...,n,$ (4)
@@
We are presented with a set of samples, each comprising a series of values $(p_{k1},p_{k2},...,p_{kn})$, referred to as the arguments, and a corresponding single value named the aggregated value, denoted as $d_{k}$. Our aim is to develop a model that accurately captures this aggregation process using the OWA method. In essence, the task is to derive a weighting vector $W$ that effectively represents the aggregation mechanism across the dataset. The aim is to determine the weighting vector $W$ based on the data provided.
@@
Minimize the immediate errors $\mathbf{e_{k}}$ :
@@
$\beta_{i}(l+1)=\beta_{i}(l)-\alpha w_{i}(b_{ki}-\hat{d}_{k})(\hat{d}_{k}-d_{k}),$ (12)
@@
$w_{i}=\frac{e^{\beta_{i}(l)}}{\sum\limits_{j=1}^{n}e^{\beta_{j}(l)}},\ \ i=[1,n],$ (13)
@@
In summary, both the DOWA and IOWA operators offer a systematic approach to aggregating probabilistic outputs by considering the order and relative similarity of predictions. By learning appropriate weights from training data and then applying them to sorted inputs, these operators capture subtle variations in classifier outputs, thereby enhancing the overall ensemble’s reliability. In the following section, we detail how these aggregation techniques are integrated into our proposed fraud detection framework, illustrating their role in the model’s multi-layer architecture.
II-D CCF detection Framework: Structure, Development, and Examination
@@
Fig. 2 illustrates the structure of the traditional stacking classifier, an ensemble learning technique that improves predictive $A_{c}$ by combining multiple ML models. In this method, several learners are trained on the dataset, and their predictions are then used as input features for a meta-learner. The meta-learner learns how to optimally combine the outputs of the first layer models to make the final prediction.